### v2

@@ 4.3 Stage II: Test Time Adaptation
Based on the insights from previous work (Wang et al, 2021), high entropy samples should be filtered out, as lower entropy typically indicates higher accuracy. Consequently, we sort the entropy of all samples in the mini-batch and select the top $\alpha\%$ samples with lower entropy, donated as $\hat{x}=\{\hat{x}_{1},\hat{x}_{2},\ldots,\hat{x}_{\lfloor\alpha\cdot B\rfloor}\}$.
@@
Table 2: Table 2: Comparison with test-time adaptation methods on three datasets with ResNet-50 backbone. FS-TTA achieves consistent improvements over TSD (Wang et al, 2021), the strongest baseline method.
@@
ERM (Vapnik, 1999) 60.7 55.7 76.2 76.8 67.4 82.5 80.8 94.0 80.9 84.5 45.2
BN (Nado et al, 2020) 58.2 55.6 75.1 75.5 66.1 83.2 84.9 94.0 77.9 85.0 43.3
TENT (Wang et al, 2021) 60.6 58.7 76.5 76.1 68.0 85.2 86.7 94.9 82.9 87.4 44.7
T3A (Iwasawa and Matsuo, 2021) 61.2 56.7 78.0 77.3 68.3 84.0 82.3 95.0 82.7 86.0 46.1
ETA (Niu et al, 2022) 58.4 55.8 75.2 75.5 66.2 83.2 84.9 94.0 77.9 85.0 46.1
LAME (Boudiaf et al, 2022) 58.7 55.6 75.1 75.4 66.2 84.9 85.5 95.0 80.9 86.6 43.2
TSD (Wang et al, 2023) 62.3 57.5 77.5 77.5 68.7 87.6 88.7 96.1 85.0 89.4 47.7
PROGRAM (Sun et al, 2024) 63.4 54.3 77.2 77.2 68.0 87.2 84.1 96.9 76.4 86.2 43.3
DEYO (Lee et al, 2024) 63.8 54.9 76.4 77.3 68.1 88.4 85.2 97.1 82.3 88.2 42.5
@@
FT+TENT (Wang et al, 2021) 68.8 65.5 79.8 78.5 73.2 87.0 86.9 95.2 83.6 88.2 45.4
FT+TSD (Wang et al, 2023) 70.5 65.1 80.3 79.2 73.8 88.3 88.6 96.5 85.9 89.8 48.5
@@
$\Delta_{up}$ over TSD (+10.9$\uparrow$ (+10.8$\uparrow$ (+5.5$\uparrow$ (+4.1$\uparrow$ (+7.8$\uparrow$ (+2.8$\uparrow$ (+1.0$\uparrow$ (+1.5$\uparrow$ (+2.8$\uparrow$ (+2.0$\uparrow$ (+3.9$\uparrow$
$\mathcal{M}_{j}=\mathbbm{1}[\arg\max p_{j}=\arg\max\hat{p}_{j}].$ (13)
5 Experiment
5.1 Experimental Settings
@@
  • PACS (Li et al, 2017) consists of 9,991 images spanning four distinct domains: Art, Cartoon (Cart), Photo (Phot), and Sketch (Sket). Each domain contains seven object categories: dog, elephant, giraffe, guitar, horse, house, and person.
  • Office-Home (Venkateswara et al, 2017) comprises 15,588 images distributed across four domains: Art, Clipart (Clip), Product (Prod), and Real-World (Real), with each domain encompassing 65 image categories.
  • DomainNet (Peng et al, 2019) is a large-scale dataset containing six domains: Clipart (Clip), Infograph (Info), Painting (Pain), Quickdraw (Quic), Real, Sketch (Sket), comprising a total of 586,575 images across 345 classes.
Implementation Details. In our main experiments, we employ ResNet-50(He et al, 2016), pre-trained on ImageNet-1k(Russakovsky et al, 2015), as the backbone model, as it is widely adopted in the test-time adaptation literature. For source model training, we follow the leave-one-domain-out protocol, as recommended by prior studies (Wang et al, 2023; Zhou et al, 2021), treating one domain as the unlabeled target and the rest as source domains.We set the batch size to 32 for each source domain and use a learning rate of 5e-5. Both the dropout probability and weight decay are set to zero. The source model is trained for 5,000 iterations, except for DomainNet, where we extend training to 15,000 iterations, following the methodology in (Cha et al, 2021). All images are resized to 224 × 224, and data augmentation is applied during source domain training, including random cropping, horizontal flipping, color jittering, and intensity adjustments. For few-shot test time adaptation, we also employ the Adam optimizer and set the batch size to The few-shot support set typically selects 5 to 16 samples per class, depending on the difficulty of the target. We carry out all experiments on NVIDIA V100 GPUs.
Table 3: Compared with existing DG and SFDA methods on OfficeHome and DomainNet.
@@
ERM (Vapnik, 1999) 60.7 55.7 76.2 76.8 67.4 64.8 22.1 51.8 13.8 64.7 54.0 45.2
DNA (Chu et al, 2022) 67.7 57.7 78.9 80.5 71.2 66.1 23.0 54.6 16.7 65.8 56.8 47.2
PCL (Yao et al, 2022) 67.3 59.9 78.7 80.7 71.6 67.9 24.3 55.3 15.7 66.6 56.4 47.7
SWAD (Cha et al, 2021) 66.1 57.7 78.4 80.2 70.6 66.1 22.4 53.6 16.3 65.5 56.2 46.7
@@
SWAD + FS-TTA 77.4 71.1 86.4 84.2 79.8 - - - - - - -
Baselines. We compare our method with various test-time adaptation (TTA) approaches, including BN(Nado et al, 2020), TENT(Wang et al, 2021), ETA(Niu et al, 2022), T3A(Iwasawa and Matsuo, 2021), LAME(Boudiaf et al, 2022), TSD(Wang et al, 2023),PROGRAM (Sun et al, 2024) and DEYO (Lee et al, 2024). Additionally, we establish new baselines by integrating fine-tuning with existing TTA methods to ensure a more comprehensive comparison. Furthermore, we compare our approach with selected methods from domain generalization, source-free domain adaptation, including DNA(Chu et al, 2022), PCL(Yao et al, 2022), SWAD(Cha et al, 2021), and F-mix(Kundu et al, 2022). Finally, we set up a comparison with the few-shot transfer learning methods, including AdaBN (Li et al, 2016), $L^{2}$ (Li et al, 2018a), DELTA (Li et al, 2019), FLUTE (Triantafillou et al, 2021), LCCS (Zhang et al, 2022). For a global overview, we compare our method with state-of-the-art method in various settings, as shown in Figure 4.
5.2 Performance Comparisons
Comparison with TTA methods. Table 2 details the comparison results between our method and various TTA methods on the Office-Home and PACS datasets, as well as the final results of DomainNet (detailed in Table 3). We observe that our method achieves state-of-the-art performance.
@@ 5.2 Performance Comparisons
Table 4: Compared with few-shot transfer learning methods on PACS dataset.
@@
AdaBN (Li et al, 2016) 85.0 83.5 96.0 78.7 85.8
$L^{2}$ (Li et al, 2018a) 85.6 84.1 96.4 76.3 85.6
DELTA (Li et al, 2019) 85.6 83.8 96.5 76.3 85.6
FLUTE (Triantafillou et al, 2021) 87.2 86.1 97.2 81.7 88.1
LCCS (Zhang et al, 2022) 87.7 86.9 97.5 83.0 88.8
@@
Finally, for a more comprehensive comparison with TTA methods, we construct new baselines by combining fine-tuning with representative TTA approaches. Specifically, we select TENT (Wang et al, 2021), as a widely adopted and foundational method in test-time adaptation, and TSD (Wang et al, 2023), which demonstrates state-of-the-art performance across benchmarks. According to the results in Table 2, our method achieves an average improvement of 4.2% over Fine-Tuning+TENT and 2.4% over Fine-Tuning+TSD across the three datasets. These results highlight the superiority of our framework in migrating to the few-shot TTA setting, benefiting from the proposed FDA module and the support-set-based prototype initialization.
Figure 5: Ablation Study on (a) effectiveness analysis about two-stage framework and (b) effectiveness analysis about FDA module.
Figure 6: Comprehensive Sensitivity Analysis of (1) Different Entropy Filter Proportion $\alpha$ and K-shot Number on Model Performance.
@@
To answer this question, we compare our method with several methods in DG and SFDA. The results of Office-Home dataset are shown in Table 3. It can be seen that our method outperforms the state-of-the-art methods in DG, such as SWAD and PCL. Furthermore, equipped with SWAD (SWAD+FS-TTA), our method achieves 79.8% accuracy. This result benefits from our adaptation of the model during the test time. In comparison to advanced SFDA methods, FS-TTA still achieves satisfactory results. It is worth noting that FS-TTA is more flexible in real-world scenarios than SFDA since it adapts the target data in an offline manner, requiring more training loops and resources. The results of DomainNet are shown in Table 3. The overall performance of FS-TTA outperforms the SFDA methods, suggesting that FS-TTA is more adept at handling challenging tasks.
Comparison with few-shot transfer learning methods. In our research, we focus on Few-Shot Test Time Adaptation (FS-TTA), which utilizes a small number of target domain samples to enhance adaptation. To ensure a comprehensive evaluation, we compare our approach with existing few-shot transfer learning methods. The results on the PACS dataset are presented in Table 4. According to the results, FS-TTA consistently outperforms all baseline methods across different domains, achieving the highest average accuracy of 91.4%, which surpasses the best-performing baseline LCCS (88.8%). This result highlights the effectiveness of our approach in adapting to domain shifts and improving classification performance in the few-shot setting.
5.3 Ablation Study
@@
Figure 7: The t-SNE feature visualization of (a) ERM, (b) TENT, (c) LAME, and (d) FS-TTA (Ours).
Figure 8: Analyzing the Impact of Batch Size on Accuracy and Running Efficiency.
@@
Sensitivity to $\alpha$. The parameter $\alpha$ represents the proportion of each batch that is selected through an entropy filter to update the prototype memory bank and the model. To evaluate the impact of $\alpha$, we conduct an experimental analysis on the Office-Home dataset by assigning $\alpha$ to 0, 0.3, 0.6, and 1, respectively. The results, as shown in Figure 6(b), demonstrate that $\alpha>0$ yields performance improvements compared to $\alpha=0$ (the source model), highlighting the effectiveness of our proposed framework. Furthermore, $\alpha=0.3$ and $\alpha=0.6$ perform better than $\alpha=1$ (no filter), indicating the effectiveness of our entropy filter strategy.
Ablation experiments on shot size. To elucidate the impact of the number of k-shots on our method, we carry out additional ablation experiments within the Office-Home dataset. The findings, illustrated in Figure 6, indicate a significant performance enhancement when the shot size ranges from 1 to 10, demonstrating a rapid performance ascension in this few-shot regime. Remarkably, even minimal shot sizes such as 1-shot and 3-shot exhibit substantial effectiveness. For instance, the 3-shot configuration achieves a 3.8% performance improvement over the TSD.
Efficiency analysis. In our main experiments, we opt for a mini-batch size of 64. To examine the variations in performance and computational efficiency with different batch size during test-time adaptation, we conduct a series of analytical experiments. As shown in Figure 8(a), we observe that accuracy experiences a gradual increase as the batch size incrementally grows, reaching a plateau around a batch size of 64. In contrast, as shown in Figure 8(b), running time exhibits a decreasing trend as the batch size grows. However, beyond a batch size of 64, the running time appears to stabilize. Consequently, for real-world applications aiming to achieve a trade-off between accuracy and computational efficiency, we suggest a batch size in the vicinity of 64.
Table 5: Comparison of our method with the baseline TSD on both ResNet and ViT-B/16 backbones across the PACS and Office-Home datasets.
@@
Scalability on Vision Transformer. We conduct experiments to verify whether our method can be applied to other architectures, such as Vision Transformer (ViT) (Dosovitskiy et al, 2021). Specifically, we adopt ViT-B/16 as the backbone and compare the baseline TSD with our approach. The results, shown in Table 5, demonstrate that our method achieves consistent improvements over TSD. On the PACS dataset, our method improves the accuracy from 90.20% (TSD) to 91.89%, while on Office-Home, it further boosts performance from 81.80% to 87.32%. These gains highlight that our approach is not limited to convolutional networks but can also optimize transformer-based architectures, making it a versatile solution for various backbone choices.
6 Conclusion

### v4

@@ IV-C Stage II: Test Time Adaptation
Based on the insights from previous work [wang2020tent], high entropy samples should be filtered out, as lower entropy typically indicates higher accuracy. Consequently, we sort the entropy of all samples in the mini-batch and select the top $\alpha\%$ samples with lower entropy, donated as $\hat{x}=\{\hat{x}_{1},\hat{x}_{2},\ldots,\hat{x}_{\lfloor\alpha\cdot B\rfloor}\}$.
@@
$\mathcal{M}_{j}=\mathds{1}[\arg\max p_{j}=\arg\max\hat{p}_{j}].$ (13)
TABLE II: Comparison with test-time adaptation methods on three datasets with ResNet-50 backbone. FS-TTA achieves consistent improvements over TSD [wang2020tent], the strongest baseline method.
@@
ERM [vapnik1999overview] 60.7 55.7 76.2 76.8 67.4 82.5 80.8 94.0 80.9 84.5 45.2
BN [nado2020evaluating] 58.2 55.6 75.1 75.5 66.1 83.2 84.9 94.0 77.9 85.0 43.3
TENT [wang2020tent] 60.6 58.7 76.5 76.1 68.0 85.2 86.7 94.9 82.9 87.4 44.7
T3A [IwasawaM21] 61.2 56.7 78.0 77.3 68.3 84.0 82.3 95.0 82.7 86.0 46.1
ETA [niu2022efficient] 58.4 55.8 75.2 75.5 66.2 83.2 84.9 94.0 77.9 85.0 46.1
LAME [boudiaf2022parameter] 58.7 55.6 75.1 75.4 66.2 84.9 85.5 95.0 80.9 86.6 43.2
PROGRAM [sun2024program] 63.4 54.3 77.2 77.2 68.0 87.2 84.1 96.9 76.4 86.2 43.3
DEYO [lee2024entropy] 63.8 54.9 76.4 77.3 68.1 88.4 85.2 97.1 82.3 88.2 42.5
TSD [wang2023feature] 62.3 57.5 77.5 77.5 68.7 87.6 88.7 96.1 85.0 89.4 47.7
@@
FT+TENT [wang2020tent] 68.8 65.5 79.8 78.5 73.2 87.0 86.9 95.2 83.6 88.2 45.4
FT+TSD [wang2023feature] 70.5 65.1 80.3 79.2 73.8 88.3 88.6 96.5 85.9 89.8 48.5
@@
[…]
V Experiment
V-A Experimental Settings
@@
  • PACS [li2017deeper] consists of 9,991 images spanning four distinct domains: Art, Cartoon (Cart), Photo (Phot), and Sketch (Sket). Each domain contains seven object categories: dog, elephant, giraffe, guitar, horse, house, and person.
  • Office-Home [venkateswara2017deep] comprises 15,588 images distributed across four domains: Art, Clipart (Clip), Product (Prod), and Real-World (Real), with each domain encompassing 65 image categories.
  • DomainNet [peng2019moment] is a large-scale dataset containing six domains: Clipart (Clip), Infograph (Info), Painting (Pain), Quickdraw (Quic), Real, Sketch (Sket), comprising a total of 586,575 images across 345 classes.
Implementation Details. In our main experiments, we employ ResNet-50[he2016deep], pre-trained on ImageNet-1k[russakovsky2015imagenet], as the backbone model, as it is widely adopted in the test-time adaptation literature. For source model training, we follow the leave-one-domain-out protocol, as recommended by prior studies [wang2023feature, zhou2021domain], treating one domain as the unlabeled target and the rest as source domains.We set the batch size to 32 for each source domain and use a learning rate of 5e-5. Both the dropout probability and weight decay are set to zero. The source model is trained for 5,000 iterations, except for DomainNet, where we extend training to 15,000 iterations, following the methodology in [swad]. All images are resized to 224 × 224, and data augmentation is applied during source domain training, including random cropping, horizontal flipping, color jittering, and intensity adjustments. For few-shot test time adaptation, we also employ the Adam optimizer and set the batch size to The few-shot support set typically selects 5 to 16 samples per class, depending on the difficulty of the target. We carry out all experiments on NVIDIA V100 GPUs.
TABLE III: Compared with existing DG and SFDA methods on OfficeHome and DomainNet.
@@
ERM [vapnik1999overview] 60.7 55.7 76.2 76.8 67.4 64.8 22.1 51.8 13.8 64.7 54.0 45.2
DNA [chu2022dna] 67.7 57.7 78.9 80.5 71.2 66.1 23.0 54.6 16.7 65.8 56.8 47.2
PCL [yao2022pcl] 67.3 59.9 78.7 80.7 71.6 67.9 24.3 55.3 15.7 66.6 56.4 47.7
SWAD [swad] 66.1 57.7 78.4 80.2 70.6 66.1 22.4 53.6 16.3 65.5 56.2 46.7
@@
SWAD + FS-TTA 77.4 71.1 86.4 84.2 79.8 69.2 31.5 57.9 25.1 70.8 62.0 52.8
Baselines. We compare our method with various test-time adaptation (TTA) approaches, including BN[nado2020evaluating], TENT[wang2020tent], ETA[niu2022efficient], T3A[IwasawaM21], LAME[boudiaf2022parameter], TSD[wang2023feature],PROGRAM [sun2024program] and DEYO [lee2024entropy]. Additionally, we establish new baselines by integrating fine-tuning with existing TTA methods to ensure a more comprehensive comparison. Furthermore, we compare our approach with selected methods from domain generalization, source-free domain adaptation, including DNA[chu2022dna], PCL[yao2022pcl], SWAD[swad], and F-mix[KunduKBMKJR22]. Finally, we set up a comparison with the few-shot transfer learning methods, including AdaBN [li2016AdaBN], $L^{2}$ [li2018explicitTL], DELTA [Li2019DELTADL], FLUTE [triantafillou2021flute], LCCS [zhang2022few]. For a global overview, we compare our method with state-of-the-art method in various settings, as shown in Figure 4.
V-B Performance Comparisons
Comparison with TTA methods. Table II details the comparison results between our method and various TTA methods on the Office-Home and PACS datasets, as well as the final results of DomainNet (detailed in Table III). We observe that our method achieves state-of-the-art performance.
@@ V-B Performance Comparisons
TABLE IV: Compared with few-shot transfer learning methods on PACS dataset.
@@
AdaBN [li2016AdaBN] 85.0 83.5 96.0 78.7 85.8
$L^{2}$ [li2018explicitTL] 85.6 84.1 96.4 76.3 85.6
DELTA [Li2019DELTADL] 85.6 83.8 96.5 76.3 85.6
FLUTE [triantafillou2021flute] 87.2 86.1 97.2 81.7 88.1
LCCS [zhang2022few] 87.7 86.9 97.5 83.0 88.8
@@
Finally, for a more comprehensive comparison with TTA methods, we construct new baselines by combining fine-tuning with representative TTA approaches. Specifically, we select TENT [wang2020tent], as a widely adopted and foundational method in test-time adaptation, and TSD [wang2023feature], which demonstrates state-of-the-art performance across benchmarks. According to the results in Table II, our method achieves an average improvement of 4.2% over Fine-Tuning+TENT and 2.4% over Fine-Tuning+TSD across the three datasets. These results highlight the superiority of our framework in migrating to the few-shot TTA setting, benefiting from the proposed FDA module and the support-set-based prototype initialization.
Figure 5: (a) Effectiveness of the two-stage framework. Stage I already yields notable gains over the source model, while adding Stage II brings further improvements, demonstrating the complementary roles of boundary alignment and distribution refinement. (b) Effectiveness of the FDA module. FDA improves few-shot fine-tuning over both the baseline and Mix-up, showing that enhanced feature diversity benefits low-shot adaptation.
@@
To answer this question, we compare our method with several methods in DG and SFDA. The results of Office-Home dataset are shown in Table III. It can be seen that our method outperforms the state-of-the-art methods in DG, such as SWAD and PCL. In addition, we combine FS-TTA with a SWAD-pretrained model to examine whether our framework can further benefit from a stronger DG-trained initialization. The resulting SWAD+FS-TTA model achieves 79.8% accuracy on OfficeHome, indicating that FS-TTA can effectively leverage the flat-minima representations learned by SWAD and yield additional improvements during test-time adaptation. This demonstrates that FS-TTA is highly compatible with diverse source-model training paradigms, serving as a plug-and-play adaptation module that can be seamlessly integrated with various DG optimizers to further enhance robustness under domain shift.In comparison to advanced SFDA methods, FS-TTA still achieves satisfactory results. It is worth noting that FS-TTA is more flexible in real-world scenarios than SFDA since it adapts the target data in an offline manner, requiring more training loops and resources. The results of DomainNet are shown in Table III. The overall performance of FS-TTA outperforms the SFDA methods, suggesting that FS-TTA is more adept at handling challenging tasks.
Comparison with few-shot transfer learning methods. In our research, we focus on Few-Shot Test Time Adaptation (FS-TTA), which utilizes a small number of target domain samples to enhance adaptation. To ensure a comprehensive evaluation, we compare our approach with existing few-shot transfer learning methods. The results on the PACS dataset are presented in Table IV. According to the results, FS-TTA consistently outperforms all baseline methods across different domains, achieving the highest average accuracy of 91.4%, which surpasses the best-performing baseline LCCS (88.8%). This result highlights the effectiveness of our approach in adapting to domain shifts and improving classification performance in the few-shot setting.
V-C Ablation Study
@@
Figure 6: (a) Sensitivity of pseudo-label filtering to different entropy-filter proportions $\alpha$. (b) Effect of K-shot supervision strength on adaptation accuracy. (c) Influence of batch size on accuracy during online refinement. (d) Influence of batch size on running efficiency.
Figure 7: t-SNE visualization of target-domain features under different adaptation methods. The plots compare the feature distributions produced by (a) ERM, (b) TENT, (c) LAME, and (d) our FS-TTA framework, illustrating how various strategies influence the target-domain representation structure.
@@
Sensitivity to $\alpha$. The parameter $\alpha$ represents the proportion of each batch that is selected through an entropy filter to update the prototype memory bank and the model. To evaluate the impact of $\alpha$, we conduct an experimental analysis on the Office-Home dataset by assigning $\alpha$ to 0, 0.3, 0.6, and 1, respectively. The results, as shown in Figure 6(a), demonstrate that $\alpha>0$ yields performance improvements compared to $\alpha=0$ (the source model), highlighting the effectiveness of our proposed framework. Furthermore, $\alpha=0.3$ and $\alpha=0.6$ perform better than $\alpha=1$ (no filter), indicating the effectiveness of our entropy filter strategy.
Ablation experiments on shot size. To elucidate the impact of the number of k-shots on our method, we carry out additional ablation experiments within the Office-Home dataset. The findings, illustrated in Figure 6(b), indicate a significant performance enhancement when the shot size ranges from 1 to 10, demonstrating a rapid performance ascension in this few-shot regime. Remarkably, even minimal shot sizes such as 1-shot and 3-shot exhibit substantial effectiveness. For instance, the 3-shot configuration achieves a 3.8% performance improvement over the TSD.
Efficiency analysis. In our main experiments, we opt for a mini-batch size of 64. To examine the variations in performance and computational efficiency with different batch size during test-time adaptation, we conduct a series of analytical experiments. As shown in Figure 6(c), we observe that accuracy experiences a gradual increase as the batch size incrementally grows, reaching a plateau around a batch size of 64. In contrast, as shown in Figure 6(d), running time exhibits a decreasing trend as the batch size grows. However, beyond a batch size of 64, the running time appears to stabilize. Consequently, for real-world applications aiming to achieve a trade-off between accuracy and computational efficiency, we suggest a batch size in the vicinity of 64.
TABLE V: Comparison of our method with the baseline TSD on both ResNet and ViT-B/16 backbones across the PACS and Office-Home datasets.
@@
Scalability on Vision Transformer. We conduct experiments to verify whether our method can be applied to other architectures, such as Vision Transformer (ViT) [dosovitskiy2020image]. Specifically, we adopt ViT-B/16 as the backbone and compare the baseline TSD with our approach. The results, shown in Table V, demonstrate that our method achieves consistent improvements over TSD. On the PACS dataset, our method improves the accuracy from 90.20% (TSD) to 91.89%, while on Office-Home, it further boosts performance from 81.80% to 87.32%. These gains highlight that our approach is not limited to convolutional networks but can also optimize transformer-based architectures, making it a versatile solution for various backbone choices.
VI Conclusion